Find all values of lemth€C such that lemth(1+2i, 5+4i) =(3+2i, 6-i)


1
Expert's answer
2021-05-11T17:34:53-0400

if "lemth \in C,lemth=a+bi"

(a+bi)(1+2i)=(a-2b)+(2a+b)i=3+2i

a=1.4,b=-0.8.

(1.4-0.8i)(5+4i)=(7+3.2)+(5.6-4)i=10.2+1.6i not equal 6-i

answer:empty set